OpenStudy (anonymous):

Multiply: 5 the square root of 8 • 4 the square root of 7 16 the square root of 10 20 the square root of 15 40 the square root of 14 40

OpenStudy (anonymous):

Under the square root, you'll have 8 x 7, right? Think of that 8 x 7 as 2*2*2*7. Is there anything you can now pull out of the square root?

OpenStudy (anonymous):

20 the square root of 15

OpenStudy (anonymous):

daddyscowgrl......it's against the website rules to just throw out answers. please don't do that. thanks @daddyscowgirl

OpenStudy (anonymous):

can you explain how you got the answer

OpenStudy (anonymous):

that answer was wrong, anyhow.

OpenStudy (anonymous):

that is the wrong answer.

OpenStudy (anonymous):

i multiplied 5 and 4 and got 20 and i added the exponents 8 and 7 and got 15

OpenStudy (anonymous):

You don't add the 8 and 7. You multiply them.

OpenStudy (anonymous):

\[5\sqrt{8}\times4\sqrt{7}=20\sqrt{8\times7}=20\sqrt{2\times2\times2\times7}=20\times2\sqrt{2\times7}=40\sqrt{14}\]

OpenStudy (anonymous):

You have 2 2's under the square root, so you can pull out a 2. Whenever you have a pair of the same factors under the square root, you can cross them off and put one of them in front of the square root

OpenStudy (anonymous):

thank you!